# Expressivity of Neural Networks with Random Weights and Learned Biases

### v1

@@ Appendix A Random Neural Network Formulation
In what follows W.L.O.G. we assume $\epsilon<1$. It is well known that we can arbitrarily approximate this dynamical system with an RNN Schäfer and Zimmermann [2006]; we provide a simple proof of this in Proposition 3. In particular, for arbitrary $01>\epsilon>$ 0 we can find an RNN of the form in Eq.2, with hidden layer width $n\in\mathbb{N}$ and output $\hat{y}$, satisfying:
$\sum_{t=1}^{T}||\hat{y}_{t}-y_{t}||<\frac{\epsilon}{2},$ (C.2)
for any initial condition of the dynamical system selected within the invariant set $U_{z}$. We note that this is a point-wise convergence. It can be shown (see Prop.3) that the hidden states of this RNN remain on a compact set, $U_{r}$, when approximating finite time trajectories of the original dynamical system with initial conditions in $U_{z}$. Given this compactness, we can then show that the following set
[…]
$r\in U_{r},x\in U,b\in[-\bar{\gamma},\bar{\gamma}],r^{\prime}\in\mathbb{R}^{n},||r^{\prime}||\leq 1\},$ (C.4)
is itself compact by the compactness of the sets from which it is formed. It will turn out that this set will contain the arguments of $\phi$ that appear in the proof. By its own compactness and the assumptions on $\phi$ we observe that $\phi$ has Lipschitz constant $K_{\phi}$. The size of $U_{r}$–and thus $\tilde{U}$–and $K_{\phi}$ will, in general, depend upon the first approximating neural network. By the compactness of $U$ and $U_{r}$, we can bound $x$ and $r$ on these sets to get bounds $R_{x}$ and $R_{r}$ respectively.
Let the parameters of the above-defined approximating RNN be given by $\theta^{*}=\{A^{*},W^{*},B^{*},b^{*}\}$. Then by Lemma 3 we can find a random RNN of hidden width $m$ and with parameters $\theta$ such that a mask, $\mathcal{M}$, exists satisfying
$|\theta^{*}_{i}-\theta_{i}^{\mathcal{M}}|<\varepsilon=\frac{\epsilon}{2lT\big{[}\max(\bar{\gamma},1)\max\big{(}n\beta K_{\phi}\tilde{R}\sum_{t=0}^{T-1}(\alpha+\bar{\gamma}\beta K_{\phi}n)^{t},1\big{)}+R_{r}\big{]}},$ (C.5)
for all $i$ with probability at least $1-\delta$, where $\tilde{R}=R_{x}+R_{r}+1$. One can show using induction (see Lemma 5 for the derivation) that if $|\theta^{*}_{i}-\theta_{i}^{\mathcal{M}}|<\varepsilon$ for all $i$ we get
$\sum_{t=1}^{T}||y_{t}^{\mathcal{M}}-\hat{y}_{t}||<lT\big{[}\bar{\gamma}n\beta K_{\phi}\tilde{R}\sum_{t=0}^{T-1}(\alpha+\bar{\gamma}\beta K_{\phi}n)^{t}+R_{r}\big{]}\varepsilon.$ (C.6)
The triangle inequality on $||y_{t}^{\mathcal{M}}-\hat{y}_{t}+\hat{y}_{t}-y_{t}||$, along with Equations C.2, C.5, and C.6 completes the proof.
@@
This follows directly from Theorem 2, by observing that one can replace the mask by simply setting biases to some sufficiently low value. ∎
Appendix D Supplementary Lemmas
The following result is well known in the literature; see e.g. Proposition 1 of Leshno et al. [1993].
@@ Appendix D Supplementary Lemmas
$=\prod_{k_{1}=1}^{k}1-P(\{k_{1}^{th}\>\mathrm{block}\>\mathrm{mask}\>\mathrm{works}\})=\bigg{[}1-\Big{(}\frac{\epsilon}{\bar{\gamma}}\Big{)}^{\pi}\bigg{]}^{\frac{m^{\star}}{n}},$ (D.2)
@@
Remark 1: We note that, in Eq.D.2, $n$ will likely also depend implicitly on $\gamma$. If $\gamma$ is very small then we will need to stack many ReLUs on top of each other to attain a large enough dynamic range to approximate the desired function (see §2.1), leading to a larger number of units. Conversely, if $\gamma$ is very large we will need to sample a large number of units before we get parameters appropriately close to the desired subnetwork configuration. This suggests the existence of some sweet spot in the value $\gamma$, which we leave for future work to explore.
Remark 2: We observe that this bound appears to be very weak. For example, if one wished to use it to find a masked network to match an MLP with input, hidden, and output dimensions of only $1$, $3$, and $1$ respectively, with a per-parameter error of $\epsilon=0.05$ an error probability of $\delta=0.1$, and with $\gamma=0.1$, this bound would suggest we need a hidden layer of $m\geq 8.34\times 10^{12}$ neurons in the bias learning network. In light of the numerical experiments, it is clear that while the math here provides proofs of existence for bias learning it does not say anything useful about the hidden layer width scaling required.
@@
For finite $0<T<\infty$, any initial condition $z_{0}\in U_{z}$, input $x_{t}\in U_{x}$ $\forall t$, $\epsilon>0$, and any $|\alpha|<\infty,\beta>0$ we can find an RNN of the style of Eq. 2 of hidden width $n\in\mathbb{N}$ and an initial value $r_{0}$ for the RNN such that:
$\sum_{t=1}^{T}||\hat{y}_{t}-y_{t}||<\epsilon$ (D.3)
where $\hat{y}_{t}$ is the output of the RNN and $y_{t}$ is that of the dynamical system. Moreover, when approximating the dynamical system in this way the RNN hidden states will remain in a compact set which we denote $U_{r}$.
The main portion of this result is well known, see e.g. Schäfer and Zimmermann [2006]. For completeness, we provide an example proof below.
@@
$z_{t+1}=F(z_{t},x_{t}),\quad y_{t}=Cz_{t},\quad z_{0}\in U_{z},$ (D.4)
defined on set $\tilde{U}_{z}=\{z_{0}+c_{0}:z_{0}\in U_{z},\>||c_{0}||<c\}$, where $U_{z}$ is an invariant set (see §2.2).
@@
$U_{zx}=\{[z+c_{0}\>\>x]:z\in U_{z},x\in U,||c_{0}||<c\}.$ (D.5)
Importantly, this set is compact given the compactness assumptions on $U$ and $U_{z}$. Also note that, since $F$ is assumed continuous, it will be $K_{F}$-Lipschitz on this compact set for some constant $K_{F}$. We can thus use the corollary to Proposition D.1 to find a neural network of hidden dimension $n\in\mathbb{N}$ that approximates $F$ with a maximum-rank output matrix, $A$. We write this neural network:
$\hat{z}=-\alpha z+\beta A\phi(Wz+Bx+b)=\hat{F}(z,x),$ (D.6)
@@
$||\hat{F}(z,x)-F(z,x)||<\varepsilon=\frac{\epsilon}{T\max(R_{C}\sum_{t=0}^{T-1}K^{t}_{F},1)},$ (D.7)
@@
$||\hat{z}_{t+1}-z_{t+1}||$ $\leq||\hat{F}(\hat{z}_{t},x_{t})-F(\hat{z}_{t},x_{t})||+||F(\hat{z}_{t},x_{t})-F(z_{t},x_{t})||$ (D.9)
$\leq\varepsilon+K_{F}||\hat{z}_{t}-z_{t}||=\varepsilon\sum_{t^{\prime}=0}^{t}K_{F}^{t^{\prime}}<\frac{c}{T}.$ (D.10)
@@
While we have approximated the dynamical system it is not yet in the standard rate-style RNN form. However, we can obtain the rate form by changing from tracking $\hat{z}$ to a different dynamical variable: $r_{t}\in\mathbb{R}^{n}$. Because $A$ is assumed to be maximum rank, we can find a minimal norm $r_{0}\in\mathbb{R}^{n}$ such that $z_{0}=Ar_{0}$, and $||r_{0}||\leq r^{\prime}$ for any $r^{\prime}$ that also satisfies $z_{0}=Ar^{\prime}$. Take this as the initial value for an RNN with dynamics:
$r_{t+1}=-\alpha r_{t}+\phi(WAr_{t}+Bx_{t}+b)$ (D.12)

### v3

@@ B.1 Random Neural Network Formulation
It is well known that we can arbitrarily approximate this dynamical system with an RNN Schäfer & Zimmermann (2006); we provide a simple proof of this in Proposition 3. In particular, for arbitrary $\epsilon>0$ we can find an RNN of the form in Eq.2, with hidden layer width $n\in\mathbb{N}$ and output $\hat{y}$, satisfying:
$\sup_{z_{0}\in U_{r},x_{0:(T-1)}\in U_{\vec{x}}}\sum_{t=1}^{T}||\hat{y}_{t}-y_{t}||<\frac{\epsilon}{2},$ (15)
Consider an $n$-width RNN of the form of Eq.2 with parameter-bounding activation functions and with initial conditions $r_{0}(z_{0})$, for $z_{0}\in U_{z}$, determined by Eq.30. This will result in $y_{t}=y_{t}(z_{0},x_{0:(t-1)},\theta)$ being a continuous function in all of its arguments. Moreover, all of its arguments are from compact sets: $U_{z}$ is compact, $U_{\vec{x}}$ is compact, and $\theta$ can be chosen to be on the hypercube with half-edges of length $\bar{\gamma}$. Thus $y_{t}$ admits a Lipschitz constant $K_{n}$. We make use of this below.
Let the parameters of the above-defined approximating RNN, from Eq.15, be given by $\theta^{*}=\{A^{*},W^{*},B^{*},b^{*}\}$. Then by Lemma 3 we can find a random RNN of hidden width $m$ and with parameters $\theta$ such that a mask, $\mathcal{M}$, exists satisfying
$|\theta^{*}_{i}-\theta_{i}^{\mathcal{M}}|<\varepsilon=\frac{\epsilon}{2K_{n}Tn(n+l+d+1)},$ (16)
for all $i$ with probability at least $1-\delta$. Then, by the Lipshitz property of $y_{t}$, we get
$\sum_{t=1}^{T}||y_{t}^{\mathcal{M}}-\hat{y}_{t}||$ $\leq\sum_{t=1}^{T}K_{n}\big{(}||\theta^{\mathcal{M}}-\theta^{*}||+||x_{0:(t-1)}-x_{0:(t-1)}||+||z_{0}-z_{0}||\big{)}$ (17)
$\leq K_{n}Tn(n+l+d+1)\varepsilon.$ (18)
This follows from $y_{t}^{\mathcal{M}}=y_{t}(r_{0}^{\mathcal{M}},x_{0:(t-1)},\theta^{\mathcal{M}})$–where we have defined $r_{0}^{\mathcal{M}}$ to be equal to ${r_{0}}_{i}$ for all $i$ such that $\mathcal{M}_{i}=1$ and 00 for all other indices–and the fact that both $r_{0}$ and $r_{0}^{\mathcal{M}}$ are continuous functions of $z_{0}$. This bound is independent of both inputs and initial condition, and, for any parameters within the hypercube defined by $\bar{\gamma}$, holds for all $z_{0}\in U_{z}$ and inputs $x_{0:(T-1)}\in U_{\vec{x}}$. By the continuity assumptions on the activation function, $f(z_{0},x_{0:(t-1)})\equiv\sum_{t=1}^{T}||y_{t}^{\mathcal{M}}-\hat{y}_{t}||$ is a continuous function of $z_{0}$ and $x_{0:(T-1)}$ for any $\theta^{*}$ and $\theta^{\mathcal{M}}$ meaning that, analogous to the feed-forward case studied previously, we can use the extreme value theorem to find maximal values $z_{0}^{\star}\in U_{z}$ and $x_{0:(t-1)}^{\star}\in U_{\vec{x}}$ such that $\sup_{z_{0}\in U_{z},x_{0:(t-1)}\in U_{\vec{x}}}f(z_{0},x_{0:(t-1)})=f(z_{0}^{\star},x_{0:(t-1)}^{\star})$. It thus follows that $\sup_{z_{0}\in U_{r},x_{0:(T-1)}\in U_{\vec{x}}}\sum_{t=1}^{T}||y^{\mathcal{M}}-\hat{y}_{t}||<\frac{\epsilon}{2}$, by Eq’s 16 and 18. The triangle inequality along with Eq.15 completes the proof.
@@
This follows directly from Lemma 2, by observing that one can replace the mask by simply setting biases to some sufficiently low value. ∎
B.4 Supplementary Lemmas
The following result is well known in the literature; see e.g. Proposition 1 of Leshno et al. (1993).
@@ B.4 Supplementary Lemmas
$=\prod_{k_{1}=1}^{k}1-P(\{k_{1}^{th}\>\mathrm{block}\>\mathrm{mask}\>\mathrm{works}\})=\bigg{[}1-\Big{(}\frac{\epsilon}{\bar{\gamma}}\Big{)}^{\pi}\bigg{]}^{\frac{m^{\star}}{n}},$ (20)
@@
Remark 1: We note that, in Eq.20, $n$ will likely also depend implicitly on $\gamma$. If $\gamma$ is very small then we will need to stack many ReLUs on top of each other to attain a large enough dynamic range to approximate the desired function (see §2.1), leading to a larger number of units. Conversely, if $\gamma$ is very large we will need to sample a large number of units before we get parameters appropriately close to the desired subnetwork configuration. This suggests the existence of some sweet spot in the value $\gamma$, which we leave for future work to explore.
Remark 2: We observe that this bound appears to be very weak. For example, if one wished to use it to find a masked network to match an MLP with input, hidden, and output dimensions of only $1$, $3$, and $1$ respectively, with a per-parameter error of $\epsilon=0.05$ an error probability of $\delta=0.1$, and with $\gamma=0.1$, this bound would suggest we need a hidden layer of $m\geq 8.34\times 10^{12}$ neurons in the bias learning network. In light of the numerical experiments, it is clear that while the math here provides proofs of existence for bias learning it massively over-estimates the layer widths required in practice, and thus does not say anything useful about the hidden layer scaling required by bias-learning.
@@
For finite $0<T<\infty$, $\epsilon>0$, and any $|\alpha|<\infty,\beta>0$ we can find an RNN of the style of Eq. 2 of hidden width $n\in\mathbb{N}$ and a continuous mapping $r_{0}:U_{z}\mapsto\mathbb{R}^{n}$ for the initial value of the RNN such that:
$\sup_{z_{0}\in U_{z},x_{0:(T-1)}\in U_{\vec{x}}}\sum_{t=1}^{T}||\hat{y}_{t}\big{(}r_{0}(z_{0}),x_{0:(T-1)}\big{)}-y_{t}(z_{0},x_{0:(T-1)})||<\epsilon$ (21)
where $\hat{y}_{t}$ is the output of the RNN and $y_{t}$ is that of the dynamical system.
The main portion of this result is well known, see e.g. Schäfer & Zimmermann (2006). For completeness, we provide an example proof below.
@@
$z_{t+1}=F(z_{t},x_{t}),\quad y_{t}=Cz_{t},\quad z_{0}\in U_{z},$ (22)
defined, by assumption, on set $\tilde{U}_{z}=\{z_{0}+c_{0}:z_{0}\in U_{z},\>||c_{0}||<c\}$, where $U_{z}$ is an invariant set (see §2.2).
@@
$U_{zx}=\{[z+c_{0}\>\>x]:z\in U_{z},x\in U,||c_{0}||<c\}.$ (23)
Importantly, this set is compact given the compactness assumptions on $U$ and $U_{z}$. Also note that, since $F$ is assumed continuous, it will be $K_{F}$-Lipschitz on this compact set for some constant $K_{F}$. By the compactness just discussed and the continuity of $F$, we can use the corollary to Proposition 19 to find a neural network of hidden dimension $n\in\mathbb{N}$ that approximates $F$ with a maximum-rank output matrix, $A$. We write this neural network:
$\hat{z}=-\alpha z+\beta A\phi(Wz+Bx+b)=\hat{F}(z,x),$ (24)
@@
$||\hat{F}(z,x)-F(z,x)||<\varepsilon=\frac{\epsilon}{T\max(R_{C}\sum_{t=0}^{T-1}K^{t}_{F},1)},$ (25)
@@
$||\hat{z}_{t+1}-z_{t+1}||$ $\leq||\hat{F}(\hat{z}_{t},x_{t})-F(\hat{z}_{t},x_{t})||+||F(\hat{z}_{t},x_{t})-F(z_{t},x_{t})||$ (27)
$\leq\varepsilon+K_{F}||\hat{z}_{t}-z_{t}||=\varepsilon\sum_{t^{\prime}=0}^{t}K_{F}^{t^{\prime}}<\frac{c}{T}.$ (28)
@@
While we have approximated the dynamical system it is not yet in the standard rate-style RNN form. However, we can obtain the rate form by changing from tracking $\hat{z}$ to a different dynamical variable, $r_{t}\in\mathbb{R}^{n}$, that satisfies $\hat{z}_{t}=Ar_{t}$. We will make a brief detour to characterize this variable.
Because $A$ is full rank, $\mathrm{col}(A)=\mathbb{R}^{s}$ so we can find an index $\nu\subset\{1,\dots,n\}$ such that $\{A_{:i}\}_{i\in\nu}$ forms a basis for $\mathbb{R}^{s}$. If we construct a matrix $A_{\nu}\in\mathbb{R}^{s\times s}$ whose columns are simply the basis vectors this matrix will have an inverse $A_{\nu}^{-1}$. We can then define the initial condition, $r_{0}$, element-wise as:
${r_{0}}_{i}(z_{0})=\begin{cases}{A_{\nu}^{-1}}_{i:}z_{0}&i\in\nu\\ 0&i\not\in\nu.\end{cases}$ (30)